OpenStudy (anonymous):

A constant force F moves an object along a straight line from the point P to the point Q. Find the work done if the distance is measured in feet and the force is measured in pounds. F = ‹6, 5› ; P = (4, 9), Q = (11, 10)

OpenStudy (anonymous):

F dot (P+Q)

OpenStudy (anonymous):

horizontal workdone=f*d for f=6p=2.72kgs;d=7ft=2.133 w.d=2.72*2.133=5.80kgs/m for f=5p=2.27kgs;d=1ft=0.3m w.d=2.27*0.3=0.681kgs/m total w.d=o.681+5.80=6.481kgs/m